OpenStudy (ciel21):

What is u + v + w when u = –10, v = 4, and w = 8? A. –22 B. 2 C. 14 D. 22

OpenStudy (anonymous):

B. Because -10 +4= -6 -6+8= 2

OpenStudy (ciel21):

thanks :P